# Collaboratively adding new knowledge to an LLM

### v1

@@ 5 Conclusion
We analyze collaboratively adding new knowledge to a large language model. In the semi-cooperative setting where datasets are not available after training, MOE mixing, model merging, and LoRA-based orthogonal subspace sequential learning perform well. It is preferable to use a small weight on the orthogonality term in the loss function to mitigate the reduction in CE loss optimization that comes from the presence of that term. MOE model mixing can mix trained LoRA adapters or FFT networks. Both perform very well but LoRA MOE Mixed models remain of modest size while FFT MOE Mixed models are limited by the considerable memory requirement as the number of experts grows. LoRA overall performs better in most cases than full-fine tuning of all parameters when both new knowledge acquisition and old, including recent, knowledge retention are taken into account along with memory requirements of the FFT MOE Mixed models.
In the fully-cooperative setting, joint training and sequential training with replay are both effective approaches with LoRA training generally preferable to full fine-tuning.
The conclusions of this work are hearttening in that LoRA has become a de-facto approach for model fine-tuning due to its considerable computational advantages combined with the plug-and-play nature of the resulting adapters. This work shows that LoRA is not only an efficient and practical solution to fine-tuning but also an effective one for collaborative and sequential training.
The high quality of MOE mixing is also notable. The method, like LoRA training, is distinguished by its computational efficiency. Given trained models the MOE can be created with little to no training using the approach provided in Lee et al. (2024). Yet, it is not only efficient but highly effective in leveraging multiple trained models or adapters.
@@ Acknowledgements
The authors would like to acknowledge Shiau Hong Lim for his deep insights and helpful discussions throughout this work.

### v2

@@ 5 Conclusion
We analyze collaboratively adding new knowledge to a large language model. In the semi-cooperative setting where datasets are not available after training, MOE mixing, model merging, and LoRA-based orthogonal subspace sequential learning perform well. It is preferable to use a small weight on the orthogonality term in the loss function to mitigate the reduction in CE loss optimization that comes from the presence of that term.
LoRA overall performs better in most cases than full-fine tuning of all parameters when both new knowledge acquisition and old, including recent, knowledge retention are taken into account. In the fully-cooperative setting, joint training and sequential training with replay are both effective approaches with LoRA training generally preferable to full fine-tuning.
The conclusions of this work are heartening in that LoRA has become a de-facto approach for model fine-tuning due to its considerable computational advantages combined with the plug-and-play nature of the resulting adapters. This work shows that LoRA is not only an efficient and practical solution to fine-tuning but also an effective one for collaborative and sequential training.
The high quality of MOE mixing is also notable. The method, like LoRA training, is distinguished by its computational efficiency. Given trained models the MOE can be created with little to no training using the approach provided in [13]. Yet, it is not only efficient but highly effective in leveraging multiple trained models or adapters.
@@ Acknowledgements
The authors would like to acknowledge Shiau Hong Lim for his deep insights and discussions throughout this work and Ronny Luss for his helpful comments on the paper.